The most recent financial statements for Heine, Inc., are shown here:

  

Income Statement Balance Sheet
  Sales $ 30,000   Assets $ 56,100   Debt $ 20,500
  Costs 22,000   Equity 35,600
  Taxable income $ 8,000   Total $ 56,100   Total $ 56,100
  Taxes (40%) 3,200
  
   Net income $ 4,800
  

  

Assets and costs are proportional to sales. Debt and equity are not. A dividend of $2,200 was paid, and the company wishes to maintain a constant payout ratio. Next year’s sales are projected to be $33,000.

  

What is the external financing needed? (Do not round intermediate calculations.)

  

  External financing needed $
Answer #1

Growth rate in sales=(33,000-30,000)/30,000=10%

Dividend payout ratio=Dividend payout/Net income

=2200/4800=0.458333333

Sales 33,000
Costs(22,000*1.1) 24200
Taxable income 8800
Taxes(8800*40%) 3520
Net income 5280
Less:Dividends(5280*0.458333333) 2420
Addition to retained earnings 2860

Total assets would be=$56100*1.1=$61710

Total equity would be=$35600+Addition to retained earnings

=35600+2860=$38460

Total assets=Total equity+Total liabilities

Hence external financing needed=$61710-38460-20500

=$2750